integral mod cosx-sinx from limits o to 3pi/2 is quetion

Grade:12th pass

1 Answers

bharat bajaj IIT Delhi
10 years ago
Integral of y = |cos x - sin x | ( 0 to 3pi/2) :
= int ( 0 to pi/4) (cos x - sin x ) + int ( pi/4 to 5pi/4) ( sin x - cos x) + int ( 5pi/4 to 3pi/2) ( cos x - sin x )
= ( 1 - v2) + (-2 v2) + ( 1 - v2)
= 2 - 4 v2
Thanks
Bharat Bajaj
IIT Delhi
askiitians faculty